# STAR-Bench: Probing Deep Spatio-Temporal Reasoning as Audio 4D Intelligence

### v1

@@ 1 Introduction
STAR-Bench is designed through a hierarchical task structure with two levels. At the Foundational Acoustic Perception level, we conduct a fine-grained, quantitative evaluation of six core audio attributes (pitch, loudness, duration, azimuth, elevation, distance) across both absolute perception ranges and relative discrimination sensitivity. We also introduce a Holistic Spatio-Temporal Reasoning level that evaluates an audio model’s ability to infer both event order and 3D scene structure. Temporal reasoning is tested via segment reordering that spans continuous processes and discrete event scripts, while spatial reasoning covers static localization, multi-source relations, and dynamic trajectory tracking. As shown in the right part of Fig. 1, every question in our holistic tasks is designed to probe a synthesis of three core pillars, such as multi-step reasoning. A failure in any one of these pillars will lead to an incorrect response. Our data curation pipeline couples procedurally synthesized, fully parameterized audio for foundational perception with large-scale real-world corpora for holistic reasoning. For the latter, we use a four-stage process including human annotation and final selection by human performance to ensure the high quality of our benchmark samples.
@@ 2 Related Work
The recent progress of Large Audio-Language Models (LALMs)(Kong et al., 2024; Chu et al., 2024; Wu et al., 2025; Xiaomi, 2025) and Omni-Language Models (OLMs)(Xu et al., 2025; Yao et al., 2024; AI et al., 2025) has significantly advanced audio understanding. At the same time, it has spurred the development of numerous benchmarks to comprehensively evaluate their capabilities. Earlier benchmarks(Wang et al., 2024; Yang et al., 2024) mainly focused on semantic-level understanding tasks (transcription, captioning, and simple question answering), and recent benchmarks(Sakshi et al., 2025; Ma et al., 2025; Kumar et al., 2025) have begun to investigate logical audio reasoning tasks. However, existing benchmarks do not address 4D audio intelligence or deep spatio-temporal reasoning across multiple audio inputs, and instead remain limited to single-clip understanding and reasoning. To fill these gaps, we propose a benchmark designed for multi-audio and deep spatio-temporal reasoning, enabling more comprehensive evaluation of audio 4D intelligence. See Table 1 for a comparison with existing benchmarks, and Appendix A for further related works.
Table 1: A comparative overview of our benchmark against other representative audio benchmarks. (: Fully supported, : Partially supported, : Not supported)
@@ 3.2.2 Spatial Reasoning Tasks
However, evaluating existing LALMs on multi-channel spatial tasks is challenging. The common practice of these models is to average multi-channel audio into a mono signal, resulting in the loss of substantial spatial information. We conduct a simple experiment as shown in Fig. 3. We construct 20 pseudo-stereo signals by assigning the original audio to the left channel and its additive inverse to the right. While human listeners could easily perform sound event classification on these signals, the models consistently failed due to signal cancellation during the mono conversion. The result confirms their lack of explicit support for genuine stereo audio processing. To provide a comprehensive assessment, we design two complementary strategies: native input, where the model directly processes stereo audio, follow their default processing pipeline to probe its intrinsic ability to exploit spatial cues; and the channel-wise input, where each channel is presented separately with explicit textual instructions, as shown in the bottom right of Fig. 2, and allow the model to approximate human-like use of interaural cues.
@@ 3.3 Data Curation Pipeline
2) AI-Assisted Automated Filtering: This process employs an efficient three-stage funnel. First, we discard unsuitable samples based on basic properties like duration and energy. Next, an LLM (e.g., DeepSeek-V3 (Liu et al., 2024a)) performs an initial screening based on textual metadata, providing justifications for its decisions. Finally, a powerful multimodal model (e.g., Gemini 2.5 Pro (Comanici et al., 2025)) analyzes the audio, metadata, and the LLM’s outputs. The final step yields a judgment, a quality score, and a preliminary classification, further filtering irrelevant samples. The detailed prompts used to query the LLMs are provided in Sec. B.3.
3) Human Annotation and Quality Control: We recruit and train 10 undergraduate annotators to label the data using a professional platform. During this process, AI-generated information is provided as an auxiliary reference. To ensure high-quality labels, we implement a stringent two-round review process: the first round involves inter-annotator cross-validation until a consensus is reached, while the second consists of random spot-checks by three domain experts.
@@ 4.1 Main Result Analysis
STAR-Bench is Challenging STAR-Bench presents a considerable challenge for existing models. Human evaluators achieve high accuracy across all task categories (e.g., 75.6% on perception, 88.0% on temporal, and 73.7% on spatial tasks), whereas all tested models fall well below this baseline. Most open-source models perform close to random guessing, and even the best closed-source model, Gemini 2.5 Pro, reaches only 49.59% average accuracy. In addition, model predictions on STAR-Bench exhibit low reliability, as evidenced by the pronounced gap between their Average Accuracy (AA) and All-Correct-Rate (ACR) scores. A detailed discussion of this issue is provided in Sec. E.1. Although the underlying audio data for the temporal tasks (e.g., FSD50K, Clotho) is commonly used for model pre-training, our novel task formulation of temporal reasoning deliberately departs from conventional audio QA formats. This design allows for a more thorough evaluation of the integrated capabilities of current models.
@@
Figure 6: Error distribution across temporal and spatial Tasks.
@@ 4.2 Discussion: Why Do Existing Models Struggle on STAR-Bench?
Figure 8: The sensitivity analysis in fine-grained perception.
@@
Lack of Human-like Sensitivity in Fine-Grained Perception. To quantify the gap in perceptual sensitivity, we present model audiograms in Fig. 9 (a)(b)(c). Fig. 9 (e)(f)(g) further track the performance of both models and human subjects on the three core acoustic attributes (pitch, loudness, and duration) as task difficulty decreases. The results reveal a stark performance gap between all models and the human baseline, particularly in the perception of fine-grained loudness differences. A clear trend is visible even for the top-performing Gemini 2.5 Pro: its accuracy, while competent on easier tasks, plummets as perceptual granularity increases. This directly corroborates our error analysis, identifying fine-grained perception as its primary bottleneck. Notably, its performance on duration perception is an exception, showcasing temporal grounding capabilities superior to those of other models by accurately assessing audio segment lengths.
@@ A.2 Audio Benchmarks
Existing audio benchmarks illustrate the rapid progress of multimodal evaluation but also expose limitations. AudioBench (Wang et al., 2024) and AIR-Bench (Yang et al., 2024) primarily focus on tasks such as automatic speech recognition (ASR), spoken question answering (SpokenQA), and audio captioning (AAC). These settings tend to reduce audio understanding to transcription or description, thereby neglecting the broader spectrum of acoustic reasoning. MMAU (Sakshi et al., 2025) and MMAR (Ma et al., 2025) further expand the scope, yet their results reveal an inherent weakness: LLMs with audio captions can achieve comparable performance to advanced LALMs, suggesting that these benchmarks probe little beyond language-level semantics. MMAU-Pro (Kumar et al., 2025) attempts to add temporal and spatial reasoning, but its scope is restricted to single-audio temporal reasoning and single static spatial reasoning.
Beyond the LALMs evaluation, multimodal benchmarks in video question answering (Cheng et al., 2025; Yang et al., 2025c) and embodied AI (Yang et al., 2025b) have emphasized temporal and spatial reasoning. However, these frameworks are predominantly grounded in the visual modality, leaving the audio modality underexplored. Real-world audio understanding frequently requires integrating information across multiple sound streams and reasoning about subtle changes in intensity, phase, or frequency—capabilities that existing benchmarks scarcely capture.
@@ B.2.1 Binaural Audio Synthesis
B.3 Prompt Used for AI-Assisted Automated Filtering of Temporal Task Data
Fig. 10 and Fig. 11 present our carefully designed prompts, which leverage LLMs and MLLMs to filter candidate data that meet the requirements of audio segment reordering.
@@ E.1 High Output Instability and Concentrated Predictions
As shown in Tab. 6, the results reveal a fundamental limitation of LALMs’ spatial understanding in perception. The native input inherently discards part of the multi-channel information during model preprocessing, which leads to a significant loss of spatial cues that are essential for fine-grained reasoning. On the other hand, the channel-wise input explicitly presents each channel with textual instructions, mitigating some of the information loss. However, as most models are not trained on multi-audio inputs, they struggle to align channel representations and to exploit interaural cues reliably. Overall, the gap between human and model performance highlights that spatial reasoning in audio remains an unsolved challenge. While channel-wise input can provide partial gains, neither strategy fully captures spatial dependencies, underscoring the need for an audio encoder that natively supports multi-channel audio input.

### v2

@@ 1 Introduction
STAR-Bench is designed through a hierarchical task structure with two levels. At the Foundational Acoustic Perception level, we conduct a fine-grained, quantitative evaluation of six core audio attributes (pitch, loudness, duration, azimuth, elevation, distance) across both absolute perception ranges and relative discrimination sensitivity. We also introduce a Holistic Spatio-Temporal Reasoning level that evaluates an audio model’s ability to infer both event order and 3D scene structure. Temporal reasoning is tested via segment reordering that spans continuous processes and discrete event scripts, while spatial reasoning covers static localization, multi-source relations, and dynamic trajectory tracking. As shown in the right part of Fig. 1, every question in our holistic tasks is designed to probe a synthesis of three core pillars, such as multi-step reasoning. A failure in any one of these pillars will lead to an incorrect response. Our data curation pipeline couples procedurally synthesized, fully parameterized audio for foundational perception with large-scale real-world corpora for holistic reasoning. For the latter, we use a four-stage process including human annotation and final selection by human performance to ensure the high quality of benchmark samples.
@@ 2 Related Work
The recent progress of Large Audio-Language Models (LALMs)(Kong et al., 2024; Chu et al., 2024; Wu et al., 2025; Xiaomi, 2025) and Omni-Language Models (OLMs)(Xu et al., 2025; Yao et al., 2024; AI et al., 2025) has significantly advanced audio understanding. At the same time, it has spurred the development of numerous benchmarks to comprehensively evaluate their capabilities. Earlier benchmarks(Wang et al., 2024; Yang et al., 2024) mainly focused on semantic-level understanding tasks (transcription, captioning, and simple question answering), and recent benchmarks(Sakshi et al., 2025; Ma et al., 2025; Kumar et al., 2025) have begun to investigate logical audio reasoning tasks.
However, existing benchmarks largely overlook audio 4D intelligence. Although some advanced benchmarks do touch upon spatio-temporal aspects, their coverage remains limited in both scale and depth. While MMAU Sakshi et al. (2025), MMAU‑Pro Kumar et al. (2025) and MMAR Ma et al. (2025) contain temporal questions, they mainly involve identifying the timing or ordering of events (e.g., when a sound occurs, which event comes first). These are primarily perceptual‑layer tasks. By contrast, our “temporal deep reasoning” tasks require understanding physical principles or causal dynamics across segments (e.g., inferring how a process evolves over time or how one event implies another), which cannot be solved by local timing cues alone. In addition, the spatial tasks in MMAR and MMAU-Pro are often restricted to single-source localization, and many items do not necessitate meaningful use of stereo cues (e.g., simple arriving vs. departing judgments). In contrast, STAR-Bench introduces a hierarchical design covering three sub-tasks in complex scenes and explicitly emphasizes stereo-cue-based reasoning.
A comparative overview of STAR-Bench and prior benchmarks is presented in Tab. 1. STAR-Bench evaluates deep spatio-temporal reasoning through tasks that go beyond surface‑level perception and instead require applying physical or causal knowledge, performing multi-step reasoning in complex real-world scenarios, and integrating information across multiple clips or events. STAR-Bench rests on a hierarchical and comprehensive task design. In addition, a rigorous data curation pipeline ensures high-quality samples, and robust evaluation strengthens the reliability.
Table 1: A comparative overview of our benchmark against other representative audio benchmarks. (: Fully supported, : Partially supported or limitted amount, : Not supported)
@@ 3.2.2 Spatial Reasoning Tasks
However, evaluating existing LALMs on multi-channel spatial tasks is challenging. The common practice of these models is to average multi-channel audio into a mono signal, resulting in the loss of substantial spatial information. We conduct a simple experiment as shown in Fig. 3. We construct 20 pseudo-stereo signals by assigning the original audio to the left channel and its additive inverse to the right. While human listeners could easily perform sound event classification on these signals, the models consistently failed due to signal cancellation during the mono conversion. The result confirms their lack of explicit support for genuine stereo audio processing. To provide a comprehensive assessment, we adopt two complementary strategies. The first is the native input setting, where the model directly processes stereo audio using its default pipeline. This allows us to probe its intrinsic ability to exploit spatial cues. The second is the channel-wise input setting, where the left and right channels are presented separately with explicit textual instructions, as shown in the bottom right of Fig. 2. This configuration serves as an ablation study to examine whether current models have any spatial capability when the binaural information is preserved at the input.
3.3 Data Curation Pipeline
@@
2) AI-Assisted Automated Filtering: This process employs an efficient three-stage funnel. First, we discard unsuitable samples based on basic properties like duration and energy. Next, an LLM (e.g., DeepSeek-V3 (Liu et al., 2024a)) performs an initial screening based on textual metadata, providing justifications for its decisions. Finally, a powerful multimodal model (e.g., Gemini 2.5 Pro (Comanici et al., 2025)) analyzes the audio, metadata, and the LLM’s outputs. The final step yields a judgment, a quality score, and a preliminary classification, further filtering irrelevant samples. The detailed prompts used to query the LLMs are provided in Sec. B.3.1.
3) Human Annotation and Quality Control: We recruit and train 10 undergraduate annotators to label the data using a professional platform. During this process, AI-generated information is provided as an auxiliary reference. To ensure high-quality labels, we implement a stringent two-round review process: the first round involves inter-annotator cross-validation until a consensus is reached, while the second consists of random spot-checks by three domain experts. More details are provided in Sec. B.3.2.
@@ 4 Evaluation
Figure 6: Error distribution across temporal and spatial Tasks.
@@ 4.1 Main Result Analysis
STAR-Bench is Challenging STAR-Bench presents a considerable challenge for existing models. Human evaluators achieve high accuracy across all task categories (e.g., 75.6% on perception, 88.0% on temporal, and 73.7% on spatial tasks), whereas all tested models fall well below this baseline. Most open-source models perform close to random guessing, and even the best closed-source model, Gemini 2.5 Pro, reaches only 49.59% average accuracy. In addition, model predictions on STAR-Bench exhibit low reliability, as evidenced by the pronounced gap between their Average Accuracy (AA) and All-Correct-Rate (ACR) scores. A detailed discussion of this issue is provided in Sec. E.1. Although the underlying audio data for the temporal tasks (e.g., FSD50K, Clotho) is commonly used for model pre-training, our novel task formulation of temporal reasoning deliberately departs from conventional audio QA formats. This design allows for a more thorough evaluation of the integrated capabilities of current models. Meanwhile, this design also serves as a diagnostic lens on the limitations of current training pipelines. The poor performance across models suggests that existing training paradigms often centered on clip‑level tagging, QA, or captioning over linguistically salient cues (e.g., using FSD50K for sound event recognition) and do not equip models with the abilities needed for audio 4D intelligence.
@@ 4.2 Discussion: Why Do Existing Models Struggle on STAR-Bench?
Figure 8: The range and sensitivity analysis in foundational perception.
@@
Lack of Human-like Range and Sensitivity in Foundational Perception. To quantify the gap in perceptual range and sensitivity, we provide detailed visualizations of model performance on our foundational perception tasks in Fig. 9. The first row of Fig. 9 presents audiograms that compare model coverage across the pitch–loudness space. Gemini 2.5 Pro achieves a much broader coverage than the other two models, where greener regions indicate higher accuracy and the covered area reflects the perceptual range. In contrast, human listeners with normal hearing are expected to achieve near-full coverage, underscoring the gap between current models and human perceptual abilities in terms of range. The second row of Fig. 9 further track the performance of both models and human subjects on the three core acoustic attributes (pitch, loudness, and duration) as task difficulty decreases. The results reveal a stark performance gap between all models and the human baseline, particularly in the perception of fine-grained loudness differences. A clear trend is visible even for the top-performing Gemini 2.5 Pro: its accuracy, while competent on easier tasks, plummets as perceptual granularity increases. This directly corroborates our error analysis, identifying fine-grained perception as its primary bottleneck. Notably, its performance on duration perception is an exception, showcasing temporal grounding capabilities superior to those of other models by accurately assessing audio segment lengths.
@@ A.2 Audio Benchmarks
Existing audio benchmarks illustrate the rapid progress of multimodal evaluation but also expose limitations. AudioBench (Wang et al., 2024) and AIR-Bench (Yang et al., 2024) primarily focus on tasks such as automatic speech recognition (ASR), spoken question answering (SpokenQA), and audio captioning (AAC). These settings tend to reduce audio understanding to transcription or description, thereby neglecting the broader spectrum of acoustic reasoning. MMAU (Sakshi et al., 2025) and MMAR (Ma et al., 2025) further extend the evaluation scope. However, their results reveal an inherent weakness—LLMs equipped with audio captions can perform on par with advanced LALMs, suggesting that such benchmarks still probe little beyond language-level semantics.
Although some advanced benchmarks, such as MMAR (Ma et al., 2025) and MMAU-Pro (Kumar et al., 2025), do touch upon spatio-temporal aspects, their coverage remains limited in both scale and depth. For instance, their temporal analysis is typically reduced to identifying the timing or order of events occurring in the audio, while spatial analysis is often limited to localizing a single sound source. In contrast, our benchmark systematically evaluates models’ temporal and spatial deep reasoning capabilities within complex, real-world physical contexts, requiring them to infer causal and dynamic relationships. Beyond audio benchmarks, multimodal benchmarks in video question answering (Cheng et al., 2025; Yang et al., 2025c) and embodied AI (Yang et al., 2025b) have emphasized temporal and spatial reasoning. However, these frameworks are predominantly grounded in the visual modality, where exploration of the audio modality remains comparatively limited. In real-world scenarios, audio understanding often depends on integrating information across multiple sound streams and reasoning about subtle changes in intensity, phase, or frequency—capabilities that existing benchmarks scarcely capture.
@@ B.2.1 Binaural Audio Synthesis
B.3 Details of the Curation Process for Reasoning Tasks
B.3.1 Prompt Used for AI-Assisted Filtering of Temporal Task Data
Fig. 10 and Fig. 11 present our carefully designed prompts, which leverage Gemini 2.5 Pro to filter candidate data that meet the requirements of audio segment reordering. Briefly, we feed the audio, its metadata, and our task description, and ask Gemini 2.5 Pro to decide, under our strict criteria of strong sequence uniqueness, semantic clarity, and high logical universality, (i) whether the audio is suitable for a reordering task, (ii) whether it reflects a continuous or discrete process, (iii) the reasoning behind its judgment, and (iv) a quality score. We adopt a conservative filtering strategy, discarding only samples explicitly marked as “not applicable”. All remaining clips, along with the model’s analysis, are then passed to professional annotators for verification and annotation. A prior LLM-based filtering step follows a similar procedure, but without audio input.
@@ B.3.1 Prompt Used for AI-Assisted Filtering of Temporal Task Data
B.3.2 Details of Human Annotation and Quality Control
Following automated filtering, each candidate sample undergoes a rigorous, multi-stage human annotation and quality control process to ensure high data quality and annotation consistency. This process is as follows:
  1. (1)
    Systematic Training: All annotators received detailed written guidelines and completed a trial annotation of 10 samples. These trials were meticulously reviewed by experts to ensure a unified understanding of the criteria.
  2. (2)
    Inter-annotator Cross-validation:
    1. (i)
      Initial Annotation: A sample is first annotated by Annotator A. The annotation content includes:
      • For Temporal Reasoning: Task compliance checks, segment boundary delineation, textual descriptions for sub-clips and the global audio, scene classification, and audio quality scoring.
      • For Spatial Reasoning: Selecting appropriate segments, task classification, and the generation of a question, the correct answer, and distractor options for the multiple-choice format.
    2. (ii)
      Review and Flagging: The annotated sample is then fully reviewed by Annotator B, who flags any inconsistencies with detailed comments and marks the sample as “failed”.
    3. (iii)
      Consensus through Negotiation: Annotators A and B then discuss all flagged issues to reach a consensus and apply corrections. During the discussions, primary sources of ambiguity are as follows:
      • For Temporal Reasoning: (a) The reasonableness of the segmented clip boundaries. (b) The existence of multiple logically plausible orderings for the segmented clips. (c) Significant discrepancies in audio quality scores.(d) Adherence to formatting and content guidelines for the captions.
      • For Spatial Reasoning: (a) Whether the spatial perception presented in the audio unambiguously aligns with the annotated answer. (b) Whether the constructed question-answer pair clearly necessitates the use of audio spatial cues for resolution. (c) Potential ambiguity in mapping the event name mentioned in the question to a specific sound in the audio. (d) The appropriate difficulty and plausibility of the distractor options.
    4. (iv)
      Expert Arbitration: In the cases where a consensus cannot be reached, the sample is escalated to an expert panel for a final decision. If the experts still cannot agree, the sample is discarded.
  3. (3)
    Expert Spot-check: After passing cross-validation, a random 10% of samples undergo a final quality check by experts to ensure consistency and accuracy. Any discovered issues are then sent back for revision.
@@ E.1 High Output Instability and Concentrated Predictions
As shown in Tab. 6, the results reveal a fundamental limitation of LALMs’ in spatial perception. The native input inherently discards part of the multi-channel information during model preprocessing, which leads to a significant loss of spatial cues that are essential for fine-grained reasoning. On the other hand, the channel-wise input explicitly presents each channel with textual instructions, mitigating some of the information loss. Despite this, most existing models are not trained to handle multi-audio inputs. As a result, they consistently struggle to align channel representations and fail to make reliable use of interaural differences. Overall, the pronounced gap between human and model performance highlights that spatial reasoning in audio remains an unsolved challenge, underscoring the need for audio encoders that natively support multi-channel audio input.